OpenStudy (anonymous):

value of 3 over square root of -64

OpenStudy (anonymous):

Irrational number 3 / sqrt (-64) = 3 / sqrt (8^2 x -1) = 3 / 8 sqrt (-1) sqrt -1 = i 3 / 8i